OpenStudy (anonymous):

Factor the following trinomial completely: x^3 - x^2 - 12x A. x(x^2 - x - 12) B. x(x + 3)(x - 4) C. x(x - 3)(x - 4) D. x(x + 3)(x + 4)

OpenStudy (nurali):

x^3 - x^2 - 12x x(x^2-x-12) x(x^2-4x+3x-12) x(x(x-4)+3(x-4)) x(x-4)(x+3)

OpenStudy (anonymous):

How about

OpenStudy (nurali):

\[\frac{ x^2+x-30 }{ x+6 }\] \[\frac{ x^2+6x-5x-30 }{ x+6 }\] \[\frac{ x(x+6)-5(x+6) }{ x+6 }\] \[\frac{ (x+6)(x-5) }{ (x+6) }\] \[(x-5)\]

OpenStudy (anonymous):

Thank you so much.

OpenStudy (nurali):

Anytime.

OpenStudy (nurali):

\[\frac{ y^2-16 }{ y-4 }\] \[\frac{ (y+4)(y-4) }{ (y-4) }\] \[(y+4)\]

OpenStudy (anonymous):

I am going to do it on a new question could you help me then?

OpenStudy (nurali):

\[\frac{ x^2-9x-36 }{ x+3 }\] \[\frac{ x^2-12x+3x-36 }{ x+3 }\] \[\frac{ x(x-12)+3(x-12) }{ x+3 }\] \[\frac{ (x-12)(x+3) }{ (x+3) }\] \[(x-12)\]